OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1

ganeshie8 (ganeshie8):

got options ?

OpenStudy (anonymous):

A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5

OpenStudy (maheshmeghwal9):

We have two roots x= -4 & x= -1 thus we have two factors (x+4) & (x+1)

OpenStudy (maheshmeghwal9):

now multiply the factors & divide the options to check the correct answer